Jack is 4 times as old as Lucy. 9 years from now the sum of their age will be 53. How old are they?

Answer by Cromlix(4381)

Hi there,
Jack is 4 times as old as Lucy.
J = 4L .........Eq(1)
9 years from now the sum of their age will be 53.
(J + 9) + (L + 9) = 53........Eq(2)
Substitute 4L from Eq(1) for J in Eq(2)
4L + 9 + L + 9 = 53
Collect like terms
4L + L = 53 - 18
5L = 35
L = 7
Lucy is 7 years old
Jack is 28 years old.
Hope this helps :-)